OpenStudy (anonymous):

Subtract, Show all of ur work be sure ur answer is simplified -3/16 - 9/16

OpenStudy (lgbasallote):

you have common denominators so it's just \[\huge \frac{-3 - 9}{16}\]

OpenStudy (anonymous):

no its -3 over top of16 then biside that its -9 over 16

OpenStudy (anonymous):

Subtract, Show all of ur work be sure ur answer is simplified

OpenStudy (lgbasallote):

\[\huge \frac{-3}{16} - \frac{9}{16} \implies \frac{-3 - 9}{16}\]

OpenStudy (lgbasallote):

common denominators is just subtract the numerators then copy the denominator

OpenStudy (anonymous):

Thats it bud Thank you very much

OpenStudy (lgbasallote):

\[\Huge \color{maroon}{\mathtt{\text{<tips hat>}}}\]